When ethane, C2H6, reacts with chlorine, Cl2, the main product is C2H5Cl, but other products containing Cl such as C2H4Cl2 are also obtained in small quantities. C2H6 + Cl2 C2H5Cl + "Other Stuff" The formation of these other products reduces the yield of C2H5Cl. Using the stoichiometry given in the chemical formula above, calculate the percent yield of C2H5Cl if the reaction of 150 g of C2H6 with 230. g of Cl2 produced 170 g of C2H5Cl.
Answer #1

Molecular weight of C2H6 is 30.07 g

Molecular weight of C2H5Cl is 64.51 g

Therefore 30.07 g of ethane will produce 64.51 g of ethyl chloride is 100 % is yield

100 % yield for 150 g of C2H6

30.07 = 64.51

150 = ?

On cross multiplying

150 x 64.51 / 30.07 = 321.8 of C2H5Cl is 100 % yield or theortical yield

Yield percentage = actual yield / theortical yield x 100

Yield percentage = 170 / 321.8 x 100 = 52.83 %